A +3nC charge is placed at (4,0,0)m and a -3nC charge is placed at (-4,0,0)m.

a) What is the electric field at (0,0,0)?

b) What is the electric field at (0,y,0) for a distance y?

c) What can you say about the direction of the electric field at any point on the x=0 plane?
Answer #1

a] Electric field is given by E= kq1/r1^2 +kq2/r2^2

= 9e9*3e-9/4^2 - 9e9*-3e-9/4^2

= 3.375 N/C

b] Here y components of the fields will cancel out, x component will add up,

Electric field E= kq1/r1^2 *4/r1 + kq2/r2^2 *4/r2

= 2*kq*4/r^3

= 2*9e9*3e-9*4/(y^2+4^2)^(3/2)

c] Direction will be parallel to x axis as y,z component will cancel out just like above case.